OpenStudy (anonymous):

verify each identity using trig. identities. 1. secxcotx = cscx 2. (1-cos^2x) csc^2=1 3. (1 + tan^2x) sin^2x = tan^2x 4. sinxcosx / (1 +cosx) (1- cosx) = cotx 5. cotx - cosx = cotx (1 - sinx)

OpenStudy (anonymous):

\[\sec x \cot x=\frac{ 1 }{ \cos x }\times \frac{ 1 }{ \tan x }=\frac{ 1 }{ \cos x }\times \frac{ \cos x }{ \sin x }=\frac{ 1 }{ \sin x }=\csc x\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

ahh o yes :D

OpenStudy (anonymous):

trying

OpenStudy (anonymous):

for the second its rule \[1-\cos ^{2}x=\sin ^{2}x\]

OpenStudy (anonymous):

then \[\sin ^{2}x \times \csc ^{2}x =1\]

OpenStudy (anonymous):

done right?

OpenStudy (anonymous):

(sin^2x)(csc^2x) is equal to 1?

OpenStudy (anonymous):

yeah because the cscx=1/sinx

OpenStudy (anonymous):

for the third one its rule \[1+\tan ^{2}x=\sec ^{2}x\]

OpenStudy (anonymous):

\[\sec ^{2}x \times \sin ^{2}x=\frac{ 1 }{ \cos ^{2}x }\times \sin ^{2}x=\tan ^{2}x\]

OpenStudy (anonymous):

5 \[\cot x-\cos x=\frac{ \cos x }{ \sin x }-\cos x=\cos x(\frac{ 1 }{ \sin x }-1)=\cos x(\frac{ 1-\sin x }{ \sin x })=\cot x(1-\sin x)\]

OpenStudy (anonymous):

4 first \[\left( 1+\cos x \times \left( 1-\cos x \right) \right)=1-\cos ^{2}x\]

OpenStudy (anonymous):

then \[ 1-\cos ^{2}x=\sin ^{2}x\]

OpenStudy (anonymous):

\[\frac{ sinxcosx }{ \sin ^{2}x }=\frac{ cosx }{ sinx }=\cot x\]

OpenStudy (anonymous):

done clear?